OpenStudy (anonymous):

just want help separating like terms in the following: xy + y' = 100x

OpenStudy (anonymous):

factor right?

myininaya (myininaya):

\[y'=x(100-y)\]

myininaya (myininaya):

divide both sides by 100-y

OpenStudy (anonymous):

ohh ok thakns

OpenStudy (anonymous):

Pro challenge: evaluate y

OpenStudy (anonymous):

\[y'=100x-xy\] \[y'=x(100-y)\] \[\frac{y'}{100-y}=x\]

OpenStudy (anonymous):

i actually don't know this, so don't trust me, trust myininaya

myininaya (myininaya):

\[\int\limits_{}^{} \frac{dy}{100-y}=\int\limits_{}^{} x dx\]

myininaya (myininaya):

it seems like you do satellie

OpenStudy (anonymous):

wow no actaully i was shooting from the hip it did say "separate" and i can read, but i know next to nothing about diffeq

OpenStudy (anonymous):

haha well thanks for the help guys. i can take it from here, but feel free to try anything you want

OpenStudy (anonymous):

let me guess, now we get \[\frac{x^2}{2}+c=\log(\text{something})\] and then exponentiate right?

myininaya (myininaya):

yep -log(something)

myininaya (myininaya):

or yeah -log(something)=log(something^(-1)) lol

myininaya (myininaya):

so either way lol

myininaya (myininaya):

something^(-1) is still something

myininaya (myininaya):

not the same something of course

OpenStudy (anonymous):

oh it is just \[\log(100-y)=\frac{x^2}{2}+c\] \[100-y=e^{\frac{x^2}{2}+c}\] how am i doing so far? then \[y=100-e^{\frac{x^2}{2}+c}\] close?

myininaya (myininaya):

aren't you missing something satellite in the first line

OpenStudy (anonymous):

I really meant that challenge for the question poster, but this works too, lols.

OpenStudy (anonymous):

could be help a brother out

OpenStudy (anonymous):

The negative, mmmyeesss

OpenStudy (anonymous):

I got the same as satellite only with a plus not a minus...

OpenStudy (anonymous):

yeah @creative, ignore me i am just messing around

OpenStudy (anonymous):

Satellites don't mess around, they orbit stuff.

OpenStudy (anonymous):

too many details to keep track of

OpenStudy (anonymous):

\[-\log(100-y)=\frac{x^2}{2}+c\] that better?

OpenStudy (anonymous):

well how long has it been since you guys took calc? probably gives enough leeway to be *allowed* to forget

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

when i took math calc hadn't been invented yet

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

When I took calc, satellites weren't invented yet.

OpenStudy (anonymous):

even natural satellites? O.o kidding :)

OpenStudy (anonymous):

just in case you're still curious satellite, this is what I did: dy/(100-y) = x dx integrate both sides to get: ln |100-y| = .5x^2 + C which after being raised to e is the same as: |100-y| = |y-100| = e^(.5x^2) * e^C = Ce^(.5x^2) so the answer: y = Ce^(.5x^2) + 100